OpenStudy (anonymous):

A town has a population of 32,000 in the year 2002; 35,200 in the year 2003; 38,720 in the year 2004; and 42,592 in the year 2005. If this pattern continues, what will the population be in the year 2015? (Points : 3) 100,430 110,473 121,520 133,672

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (the_fizicx99):

Did you determine whether it's arithmetic or geometric?

OpenStudy (anonymous):

im thinking its geometric

OpenStudy (the_fizicx99):

It is, it has a ratio of 1.1 (1.10) Use the formula for explicit, it'll say you some time. a_1 = 32,000 r = 1.1 a_13 = 32,000 \(\ \sf \times 1.1^{13} \rightarrow 32,000 \times 1.1^{13}\) a_13 = 32000 * 3.4522712143931 = 110,472.6788605792 So B

OpenStudy (the_fizicx99):

2002 was when it was 0, 2015 is the estimated date. 13 - 0 = 13